Two charges - 16 and + 4 µC, are fixed in placed and separated by 3.0 m .

a)       At what spot along a line through the charges is the net electric field zero? Locate this spot relative to the positive charge. ( Hint: The spot does not necessary lie between two charges.)

b)       What would be the force on a charge of + 14 µC placed at this point?


1
Expert's answer
2021-04-19T17:10:38-0400

a)

Consider position P from the positive charge at which net electric field is zero.

The net electric field at point P is zero. So,

"\frac{kq_1}{x^2}=\frac{kq_2}{(3.0-x)^2}"

"\frac{4 \mu C}{x^2}=\frac{16 \mu C}{(3.0-x)^2}"

On solving the above equation, we get x= 1.0 m.

So the net electric field is zero at distance x=1.0 m from the positive charge.

b)

The net electric force on a charge of + 14 µC placed at point P is zero. Because electric field is zero at this point.